Session: Feb 2006 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click

Experiment Feedback

Urumqi (Zhang Hua at Thu Feb 16 23:50:28 2006) Success :

Shanghai (Tao An at Fri Feb 17 01:40:44 2006) Success :
Weather : Cloudy and rainy. no known problem.

Onsala (Michael Lindqvist at Fri Feb 17 07:45:02 2006) Success :
Cloudy weather, some snow. Otherwise no known problems.

Torun (Kaz Borkowski at Fri Feb 17 08:05:57 2006) Success :
No problems noted, although by now we know there were no FTP fringes in channel 8 (at the station all BBCs seemed to behave quite normal throughout this experiment).

Medicina (G.Maccaferri at Fri Feb 17 08:22:02 2006) Success :

Westerbork (Tony Foley at Fri Feb 17 12:40:29 2006) Minor failures :
Wb used in phased array mode. No known problems

Noto (P.Cassaro at Fri Feb 17 22:15:42 2006) Success :

Effelsberg (dgraham at Sat Feb 18 08:08:51 2006) Minor failures : HIGH
Occasional heavy rain

Onsala (Michael Lindqvist at Mon Feb 20 08:55:34 2006) Minor failures :
Antenna tracking problem a few minutes around 23.32 UTC. Cloudy.

Hartebeesthoek (Jonathan Quick at Mon Feb 20 22:42:01 2006) Minor failures : RFI
Very strong RFI (flat passband from 121.5 to 128.5 MHz in the IF) outside the recorded IF from 15:45 to 15:52 UT.

Jodrell_Bank (Alastair Gunn at Fri Mar 3 17:02:08 2006) Success :
The MkII telescope was used (Jb2). VC06 had an LO problem and was swapped for VC13 at about 14:56 UT. No other known problems.

EVN experiment feedback facility: entire session

Entire Session

Session: Feb 2006 (session 1)
Wavelengths: 6 3.6 5 18 cm


To give feedback on this entire session, click

Entire Session Feedback

Torun (Kaz Borkowski at Mon Feb 20 11:25:29 2006) Minor failures : PCAL
Tr phase cal system failed to produce signals during all the 6 cm experiments for yet unknown reason.

Westerbork (Tony Foley at Fri Feb 24 15:48:24 2006) Minor failures : ANT POIN
The feedback from the antenna control for SINGLE-DISH observations is unreliable for 20seconds after a source change. We will try to fix this. (Affected only the 5cm experiments -Z.P.)